OpenStudy (kewlgeek555):

Choose the graph below that correctly represents the equation 3x + 9y = -18.

OpenStudy (kewlgeek555):

This is a multiple choice question, but I just need a brain freshener on how to loo for a graph that correctly represents this equation.

OpenStudy (campbell_st):

ok... so the easy way to do this question is findout where the equation cuts the x and y axes x axis: let y = 0 so you are solving 3x = -18 ordered pair (answer, 0) y:axis let x = 0 so you are solving 9y = -18 ordered par (0, answer) match the intercepts to the graphs

OpenStudy (anonymous):

3x + 9y = -18 9y= -3x -18 y= -1/3x - 2

OpenStudy (kewlgeek555):

Thank you @campbell_st and @Whiterabbit541

OpenStudy (kewlgeek555):

@Whiterabbit541 do you think you could help me check some more?

OpenStudy (anonymous):

sure

OpenStudy (kewlgeek555):

Thanks. Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown. http://orange.flvs.net/webdav/assessment_images/educator_algebra1/pictures/v8/04/12/04_12_group21_graph1.gif y = -three halvesx y = three halvesx - 6 y = -two thirdsx - 5 thirds y = two thirdsx - 13 thirds

OpenStudy (anonymous):

I'm unable to answer this question because the link you have given leads to an error page.

OpenStudy (kewlgeek555):

Oh. Sorry, well let me [try] to explain the coordinate plane. The x-intercept is -5 and the y-intercept is 3. Then, there is a period on the point (2, -3)

OpenStudy (kewlgeek555):

AND - the slope is poitive.

OpenStudy (kewlgeek555):

Actuslly, now that I think of it...I think its D. It can't be A or C because those are negative. And I don't think it would be B because, Bs slope is too much. Do you think I'm right?

OpenStudy (anonymous):

Seems right to me.

OpenStudy (kewlgeek555):

Thanks. Okay. These next two are just to check if I'm right. I hope I am not overloading you @Whiterabbit541 The functions f(x) and g(x) are described below. f(x) = 3x - 7 g(x) = 3x + 8 The graph of g(x) is obtained by shifting up the graph of f(x) by _____ units. Based on my calculations I put 15. Gina pumps 40 cubic inches of air per second into a balloon. The function below shows the total amount of air f(t) inside the balloon after t seconds. f(w) = 40t + 5 Which statement best describes the y-intercept of the graph of the function? It is equal to 40 and represents the amount of air already present in the balloon before Gina starts blowing. It is equal to 5 and represents the amount of air already present in the balloon before Gina starts blowing. It is equal to 5 and represents the amount of air in the balloon after 1 second. It is equal to 45 and represents the amount of air in the balloon after 1 second. I chose B.

OpenStudy (atlas):

and you are right

OpenStudy (kewlgeek555):

Thank you so much. Sorry, but I am going to ask one more question to you. For the function y = f(x), what is the ordered pair for the point on the graph when x = 3n - 4? (3n - 4, f(n)) (3n - 4, f(3n - 4)) (x, f(3n)) (x, 3n - 4)

OpenStudy (atlas):

and what do you think is the answer?

OpenStudy (kewlgeek555):

Well, I have no idea. I know that when you want to change a slope-thingie equation to a function, you relace the y with the f(x).

OpenStudy (atlas):

you don't need to think it that way........ Any point has two coordinates x and y it is given x = 3n-4 and y=f(x) =f(3n-4) So point P can be plotted by P(x,y)

OpenStudy (atlas):

I hope you can get to the answer now

OpenStudy (kewlgeek555):

Now I get it, so it woyuld be B?

OpenStudy (atlas):

yeah right!

OpenStudy (kewlgeek555):

Oh thank you, everything understands now.